# Correcting Negative Bias in Large Language Models through Negative Attention Score Alignment

### v1

@@ Appendix I Samples of Attention Heads
StrategyQA LLaMA3-8B 0.877 / 0.877 0.864 / 0.895 0.875 / 0.836 0.869 / 0.864 55.0 / 96.2 0.102 / 0.128
Mistral-7B 0.830 / 0.849 0.802 / 0.846 0.861 / 0.837 0.830 / 0.841 34.3 / 71.9 0.116 / 0.114
Gemma-7B 0.771 / 0.774 0.907 / 0.907 0.690 / 0.692 0.784 / 0.785 71.8 / 79.2 0.162 / 0.161
Qwen2-7B 0.844 / 0.836 0.907 / 0.915 0.742 / 0.716 0.816 / 0.804 107.2 / 194.9 0.106 / 0.115
MuSique LLaMA3-8B 0.757 / 0.741 0.815 / 0.844 0.683 / 0.609 0.743 / 0.707 64.0 / 118.6 0.039 / 0.055
Mistral-7B 0.707 / 0.704 0.802 / 0.821 0.594 / 0.558 0.682 / 0.664 110.7 / 185.9 0.212 / 0.217
Gemma-7B 0.503 / 0.510 0.834 / 0.836 0.555 / 0.552 0.667 / 0.665 96.2 / 107.1 0.276 / 0.274
Qwen2-7B 0.639 / 0.622 0.858 / 0.873 0.552 / 0.500 0.672 / 0.636 282.9 / 405.7 0.200 / 0.215
GSM8K (Rephrased) LLaMA3-8B 0.545 / 0.558 0.531 / 0.551 0.838 / 0.688 0.650 / 0.612 74.7 / 123.6 0.170 / 0.105
Mistral-7B 0.512 / 0.511 0.509 / 0.515 0.782 / 0.463 0.617 / 0.487 77.0 / 127.4 0.269 / 0.294
Gemma-7B 0.524 / 0.517 0.541 / 0.542 0.887 / 0.888 0.672 / 0.673 66.2 / 74.7 0.346 / 0.349
Qwen2-7B 0.576 / 0.560 0.614 / 0.634 0.738 / 0.606 0.670 / 0.620 82.9 / 141.6 0.115 / 0.085
MATH (Rephrased) LLaMA3-8B 0.549 / 0.572 0.534 / 0.562 0.874 / 0.740 0.663 / 0.639 60.8 / 103.6 0.175 / 0.118
Mistral-7B 0.526 / 0.533 0.522 / 0.549 0.685 / 0.384 0.592 / 0.452 109.6 / 152.8 0.260 / 0.299
Gemma-7B 0.543 / 0.539 0.553 / 0.555 0.871 / 0.867 0.677 / 0.677 64.0 / 73.0 0.340 / 0.338
Qwen2-7B 0.581 / 0.549 0.617 / 0.639 0.715 / 0.611 0.662 / 0.624 103.1 / 187.5 0.162 / 0.168
AR-LSAT (Rephrased) LLaMA3-8B 0.518 / 0.510 0.537 / 0.536 0.444 / 0.337 0.486 / 0.414 63.6 / 109.2 0.227 / 0.245
Mistral-7B 0.530 / 0.526 0.553 / 0.556 0.447 / 0.382 0.494 / 0.453 80.02 / 121.3 0.355 / 0.382
Gemma-7B 0.510 / 0.514 0.523 / 0.526 0.534 / 0.551 0.528 / 0.539 68.0 / 75.6 0.372 / 0.383
Qwen2-7B 0.496 / 0.559 0.528 / 0.643 0.182 / 0.574 0.270 / 0.607 207.3 / 204.4 0.359 / 0.384
@@
LLaMA3-8B StrategyQA 0.877 / 0.870 0.864 / 0.924 0.875 / 0.789 0.869 / 0.851 55.0 / $\mathbf{-507.6}$ 0.102 / 0.176
@@
Figure 5: Confidence histograms of LLaMA3-8B before and after NASA.
Figure 6: Examples of attention heads of LLaMA3-8B-Instruct model for a GSM8K dataset sample.
Figure 7: Examples of attention heads of LLaMA3-8B-Instruct model with NASA for a GSM8K dataset sample.

### v2

@@ Appendix I Generalization to Universal Binary Decision
StrategyQA LLaMA 0.877 / 0.877 0.864 / 0.895 0.875 / 0.836 0.869 / 0.864 55.0 / 96.2 0.102 / 0.128
Mistral 0.830 / 0.849 0.802 / 0.846 0.861 / 0.837 0.830 / 0.841 34.3 / 71.9 0.116 / 0.114
Gemma 0.771 / 0.774 0.907 / 0.907 0.690 / 0.692 0.784 / 0.785 71.8 / 79.2 0.162 / 0.161
Qwen 0.844 / 0.836 0.907 / 0.915 0.742 / 0.716 0.816 / 0.804 107.2 / 194.9 0.106 / 0.115
MuSiQue (Rephrased) LLaMA 0.757 / 0.741 0.815 / 0.844 0.683 / 0.609 0.743 / 0.707 64.0 / 118.6 0.039 / 0.055
Mistral 0.707 / 0.704 0.802 / 0.821 0.594 / 0.558 0.682 / 0.664 110.7 / 185.9 0.212 / 0.217
Gemma 0.503 / 0.510 0.834 / 0.836 0.555 / 0.552 0.667 / 0.665 96.2 / 107.1 0.276 / 0.274
Qwen 0.639 / 0.622 0.858 / 0.873 0.552 / 0.500 0.672 / 0.636 282.9 / 405.7 0.200 / 0.215
GSM8K (Rephrased) LLaMA 0.545 / 0.558 0.531 / 0.551 0.838 / 0.688 0.650 / 0.612 74.7 / 123.6 0.170 / 0.105
Mistral 0.512 / 0.511 0.509 / 0.515 0.782 / 0.463 0.617 / 0.487 77.0 / 127.4 0.269 / 0.294
Gemma 0.524 / 0.517 0.541 / 0.542 0.887 / 0.888 0.672 / 0.673 66.2 / 74.7 0.346 / 0.349
Qwen 0.576 / 0.560 0.614 / 0.634 0.738 / 0.606 0.670 / 0.620 82.9 / 141.6 0.115 / 0.085
MATH (Rephrased) LLaMA 0.549 / 0.572 0.534 / 0.562 0.874 / 0.740 0.663 / 0.639 60.8 / 103.6 0.175 / 0.118
Mistral 0.526 / 0.533 0.522 / 0.549 0.685 / 0.384 0.592 / 0.452 109.6 / 152.8 0.260 / 0.299
Gemma 0.543 / 0.539 0.553 / 0.555 0.871 / 0.867 0.677 / 0.677 64.0 / 73.0 0.340 / 0.338
Qwen 0.581 / 0.549 0.617 / 0.639 0.715 / 0.611 0.662 / 0.624 103.1 / 187.5 0.162 / 0.168
AR-LSAT (Rephrased) LLaMA 0.518 / 0.510 0.537 / 0.536 0.444 / 0.337 0.486 / 0.414 63.6 / 109.2 0.227 / 0.245
Mistral 0.530 / 0.526 0.553 / 0.556 0.447 / 0.382 0.494 / 0.453 80.02 / 121.3 0.355 / 0.382
Gemma 0.510 / 0.514 0.523 / 0.526 0.534 / 0.551 0.528 / 0.539 68.0 / 75.6 0.372 / 0.383
Qwen 0.496 / 0.559 0.528 / 0.643 0.182 / 0.574 0.270 / 0.607 207.3 / 204.4 0.359 / 0.384
@@
LLaMA StrategyQA 0.877 / 0.870 0.864 / 0.924 0.875 / 0.789 0.869 / 0.851 55.0 / $\mathbf{-507.6}$ 0.102 / 0.176
@@
Figure 5: Confidence histograms of LLaMA before and after NASA.
Figure 6: Examples of attention heads of LLaMA for a GSM8K dataset sample.
Figure 7: Examples of attention heads of LLaMA model with NASA for a GSM8K dataset sample.